# MetaMixer Is All You Need

### v1

@@ 4.4 Super-Resolution
EDSR [51] CVPRW17 1488.4 196.6 3.6 1.4 DIV2K 38.11/0.9602 33.92/0.9195 32.32/0.9013 32.93/0.9351 39.10/0.9773
SwinIR-light [50] ICCVW21 1764.3 124.9 OOM OOM 38.14/0.9611 33.86/0.9206 32.31/0.9012 32.76/0.9340 39.12/0.9783
ELAN-light [114] ECCV22 687.0 50.2 OOM OOM 38.17/0.9611 33.94/0.9207 32.30/0.9012 32.76/0.9340 39.11/0.9782
DITN [56] MM23 679.1 62.8 OOM OOM 38.17/0.9611 33.79/0.9199 32.32/0.9014 32.78/0.9343 39.21/0.9781
SwinIR-NG [13] CVPR23 1712.4 143.9 - - 38.17/0.9612 33.94/0.9205 32.31/0.9013 32.78/0.9340 39.20/0.9781
$\text{FFNet}_{\textit{sr}}$-light - 394.1 55.3 1.0 0.4 38.23/0.9613 33.75/0.9204 32.34/0.9018 32.81/0.9344 39.23/0.9781
OmniSR [95] CVPR23 892.2 76.6 - - DF2K 38.29/0.9617 34.27/0.9238 32.41/0.9026 33.30/0.9386 39.53/0.9792
EDT-T [46] IJCAI23 5364.7 334.1 - - 38.23/0.9615 33.99/0.9209 32.37/0.9021 32.98/0.9362 39.45/0.9789
DAT-light [11] ICCV23 2295.2 1936.3 OOM OOM 38.24/0.9614 34.01/0.9214 32.34/0.9019 32.89/0.9346 39.49/0.9788
SAFMN-large [84] ICCV23 767.5 124.9 OOM 0.8 38.28/0.9616 34.14/0.9220 32.39/0.9024 33.06/0.9366 39.56/0.9790
$\text{FFNet}_{\textit{sr}}$ 397.6 72.3 1.4 0.6 38.32/0.9618 34.28/0.9235 32.43/0.9029 33.23/0.9382 39.76/0.9793
4.5 3D Semantic Segmentation
@@
Architecture. Following [104], we adopt the U-Net [76] structure, comprising four stages of encoders and decoders with block depths of [2, 2, 6, 2] and [1, 1, 1, 1], respectively. We also utilize the Grid Pooling introduced in [104], setting all grid size multipliers to $\times$2, which represents the expansion ratio over the preceding pooling stage. For token mixing, we employ submanifold convolution [31]. Additionally, to manage memory-constrained scenarios requiring small batch sizes, we use both batch normalization and pre-norm (layer normalization) scheme. Due to the implementation limitations of submanifold convolution, which does not support depthwise operation, we restrict our kernel size to 3$\times$3 to handle overfitting issues. Despite these constraints, the obtained results are promising, demonstrating that even small kernels are highly effective within the U-Net structure. Detailed descriptions of architecture can be found in Appendix G.3.
Table 5: Multivariate time series long-term forecasting results with prediction lengths $S\in\{96,192,336,720\}$ and fixed lookback length $T=96$. Results are averaged from all prediction lengths. ETT results further averaged by subsets. The baseline numbers are from [57]. The best results are in bold and the second best are underlined.
@@
0.265
0.317
0.292
0.363
0.309
0.360
0.288
0.314
0.338
0.382
Table 6: 3D semantic segmentation results.
3D Sem. Seg. ScanNet [17] ScanNet200 [77] S3DIS [1] Model Val Test Val Test Area5 6-fold PointNeXt [72] 71.5 71.2 - - 70.5 74.9 ST [45] 74.3 73.7 - - 72.0 - OctFormer [97] 75.7 76.6 32.6 32.6 - - Swin3D [109] 76.4 - - - 72.5 76.9 OA-CNNs [70] 76.1 75.6 32.3 33.3 71.1 - PTv2 [104] 75.4 74.2 30.2 - 71.6 73.5 FFNet 76.1 76.5 33.8 34.5 72.6 77.1
Setup. We conduct experiments using our models on the representative datasets: ScanNetv2 [17], ScanNet200 [77], and the S3DIS [1] for indoor scenes. We use Pointcep [15] codebase specialized for point cloud perception tasks. Details on the training settings are provided in Appendix H.1.
Results. Remarkably, simply adapting our methods to point clouds demonstrates superior performance compared to task-specific competitors, as shown in Tab. 6. We anticipate that incorporating adaptive convolution [70] or large kernel convolution [10] to our mixer design could further enhance performance.
4.6 Time-Series Forecasting
@@
Architecture. For time-series forecasting, we employ a backbone-head structure with RevIN [42], where the backbone consists of stacked FFNet blocks. In our mixer design, we utilize a 51-sized kernel for depthwise convolution to effectively catch up the global receptive fields of transformer-based and MLP-based models. In fact, directly applying the FFNet block (Fig. 3 (c)) to time series data is challenging due to the presence of variable dimension in addition to channel and temporal dimensions in multivariate setting. To effectively capture both temporal and cross-variable dependencies, we modify FFN part using appropriate grouping and feature reshaping. Detailed descriptions of these time series-specific modifications and architecture details are available in Appendix G.4.
@@
Setup. We conduct long-term forecasting experiments on popular 8 real-world datasets, including ETT (4 subsets), ECL, Exchange, Traffic, Weather. Following [57], we set the input length as 96 for all the results. We compute the Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics for multivariate time series forecasting. Detailed settings are provided in Appendix H.2.
@@
Results. Tab. 5 shows the comparable performance of FFNet in multivariate time-series forecasting. Specifically, FFNet consistently outperforms recent state-of-the-art transformer-based and MLP-based models in most cases, demonstrating the excellent task-generality of our approach.
5 Conclusion
@@
In this work, we analyzed the FFN from a seldom-explored perspective and integrated self-attention and FFN-based mixers within a query-key-value framework. We further abstracted this framework into a general mixer architecture, termed MetaMixer, where sub-operations are unspecified. Based on the MetaMixer, we presented a convolutional mixer design with FFNified attention. In numerous tasks where transformer-based and convolution-based models are in fierce competition, we merge both approaches in a completely new direction, and achieve a new state of the art. This achievement underscores that the significance lies not in self-attention per se, but in satisfying task-specific functionalities, starting with MetaMixer as the foundation.
@@ Appendix A FFNs Are Key-Value Memories
We also find that $K_{548}$ and $K_{1461}$ from the smaller variant PoolFormer-S12, as well as $K_{382}$ and $K_{2869}$ from the ViT-B trained on ImageNet-21K, operate similarly to the special keys discussed above. Additionally, Li et al. [49] observed that the FFN’s activations in ViT models are sparse, mirroring observations in Fig. 2 (a). The above results suggest that FFNs generally operate as key-value memories across various models.
@@ Appendix B MetaMixer Generality
As shown in Fig. 8, the most commonly used mixers are instantiated by specifying sub-operations of the MetaMixer, underscoring that the query-key-value mechanism is a essential requirement for competitive performance. Specifically, The spatial MLP [^1] Both MLP (Multi-Layer Perceptron) and FFN (Feed-Forward Network) generally refer to modules containing two fully-connected layers with an activation function between them. In this paper, we refer to this module as FFN overall, but use MLP specifically in the context of MLP-Mixer [88]. in MLP-Mixer [88] operates via the same query-key-value mechanism as an FFN, but applies it along different axes—token and channel, respectively. The ConvNeXt [60] block can be interpreted as employing a 7$\times$7 depthwise convolution as the query projection layer within an FFN. Additionally, self-attention and FFNified attention utilize inter-token matrix multiplication and depthwise convolution, respectively, to compute query-key-value interactions.
@@ Appendix C Detailed Comparisons with Related Work
MLP-Mixer. The term “FFNification” might initially evoke thoughts of spatial MLPs [88] rather than our proposed FFNified attention due to its direct implications. However, our use of the term encompasses not only the adaptation of mechanisms but also aims to achieve the generality and efficiency characteristic of FFNs. Spatial MLPs, on the other hand, suffer from parameter inefficiency [88, 119] and are impractical for high-resolution inputs, similar to the global self-attention. They also unable to handle multiple resolutions, limiting their applicability across various tasks. Consequently, depthwise convolution emerges as an efficient and effective alternative, addressing these issues. Unlike spatial MLPs, which share the same kernel across all channels, depthwise convolution applies different kernels to each channel. To bridge the receptive field gap between convolution and spatial MLP, we employ large kernels (e.g., 7$\times$7 or 9$\times$9 for image recognition and 51 for time series).
ConvNeXt. Both the ConvNeXt [60] block and FFNified attention mixer share basic elements such as convolution and GELU activation, yet they differ in the operating axes of the query-key-value mechanism. FFNified attention actively extracts spatial features through two consecutive large-kernel convolutions, calculating query-key and coefficient-value interactions. In contrast, the ConvNeXt block uses depthwise convolution as a means of query projection, indirectly utilizing the surrounding context during channel interaction. Hence, employing our proposed FFNified attention and ConvNeXt block as token mixer and channel mixer, respectively, provides a compact mixer design space. In FFNified attention, it is natural to not expand channels, leading our FFNet to maintain a larger base channel size compared to similar-sized baseline models and thus reduce the expansion ratio in the channel mixer to 3. This decision aligns with prior research findings [55] that the expanded features of channel mixer contain substantial redundancy. Thanks to our comprehensive design strategy and compact mixer space, our models demonstrate a better performance-speed trade-off than the ConvNeXt models especially on resource-constrained devices. (see Tab. 7).
ViT and Swin Transformer. Self-attention operates at the token level, employing matrix multiplication to calculate inter-token similarity and aggregate values using the resulting coefficients. In contrast, FFNified attention operates at the channel level, detecting local patterns through depthwise convolution and applying a subsequent convolution to the resulting coefficients to aggregate values. To reduce parameter complexity, self-attention shares attention weights across channels, while FFNified attention shares them across positions. Another key difference lies in the key-value generation process: self-attention utilizes instance-specific key-value pairs through linear projections, while FFNified attention initializes key-values as static weights and updates them on the training set.
Vanilla self-attention, being a token-based module, faces limitations in vision tasks with long token lengths. Swin transformer [59] addresses this by performing self-attention within local windows. However, window attention necessitates complex reshape operations, hindering efficient deployment across various inference platforms [55, 111]. Conversely, FFNified attention, composed solely of hardware-friendly operations like convolution and GELU, exhibits high speed across diverse devices (see Tab. 7).
@@
Table 7: Comparison of inference speeds across various hardware platforms for FFNet, Swin Transformer, and ConvNeXt. “bs” denotes batch size. For CPU benchmarks, models are converted to ONNX runtime format and evaluated on an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz processor using a single thread. For Mobile and M2Air, we export the models using CoreML tools and report the median latency over 1,000 runs.
Model Top-1 GPU Throughput (img/sec) CPU Throughput (img/sec) Mobile Macbook Air (%) bs 1 bs 64 bs 1 bs 16 Latency (ms) Latency (ms) Swin-T 81.3 - 1934 30 35 24.8 6.9 ConvNeXt-T 82.1 789 2397 40 48 3.7 2.4 FFNet-2 82.9 674 2142 28 52 3.1 2.1 Swin-B 83.5 - 841 13 14 33.5 12.1 ConvNeXt-B 83.8 399 1246 16 17 8.1 5.6 FFNet-3 83.9 490 1420 20 40 4.5 2.9
Appendix D Effective Receptive Field (ERF)
@@
ConvNeXt-T 1% 1.6% 4.4% 53.2%
Swin-T 1.6% 2.9% 6.5% 60.3%
FFNet-2 10.4% 16.4% 30.2% 97.9%
@@
Figure 9: Left. Effective receptive field (ERF) comparison. FFNet has broader ERFs compared to Swin Transformer and ConvNeXt, despite using the same kernel (window) size. Right. Quantitative analysis on the ERF with the high-contribution area ratio $r$. A higher value of $r$ indicates a smoother distribution of high-contribution pixels, which implies a larger ERF.

### v2

@@ 4.1 FFNet for Efficient Image Recognition
ELAN-light [112] ECCV’22 169.6 19.0 OOM 521.4 DIV2K 32.43/0.8975 28.78/0.7858 27.69/0.7406 26.54/0.7982 30.92/0.9150
SwinIR-NG [14] CVPR’23 304.8 248.6 - - 32.44/0.8980 28.83/0.7870 27.73/0.7418 26.61/0.8010 31.09/0.9161
SRFormer-light [123] ICCV’23 312.2 214.0 OOM OOM 32.51/0.8988 28.82/0.7872 27.73/0.7422 26.67/0.8032 31.17/0.9165
MambaIR-l [32] ECCV’24 400.1 51.9 - - 32.42/0.8977 28.74/0.7847 27.68/0.7400 26.52/0.7983 30.94/0.9135
SMFANet$+$ [118] ECCV’24 72.2 12.1 OOM 653.2 32.43/0.8979 28.77/0.7849 27.70/0.7400 26.45/0.7943 31.06/0.9138
$\text{FFNet}_{\textit{sr}}$-light - 87.8 10.1 110.6 99.5 32.52/0.8993 28.88/0.7884 27.77/0.7424 26.66/0.8018 31.32/0.9170
OmniSR [96] CVPR’23 238.8 24.6 - - DF2K 32.57/0.8993 28.95/0.7898 27.81/0.7439 26.95/0.8105 31.50/0.9192
DAT-light [12] ICCV’23 642.3 530.3 OOM OOM 32.57/0.8991 28.87/0.7879 27.74/0.7428 26.64/0.8033 31.37/0.9178
SAFMN-large [84] ICCV’23 216.2 19.1 OOM 217.9 32.65/0.9005 28.96/0.7898 27.82/0.7440 26.81/0.8058 31.59/0.9192
SMFANet$+$ [118] ECCV’24 72.2 12.1 OOM 653.2 32.51/0.8985 28.87/0.7872 27.74/0.7412 26.56/0.7976 31.29/0.9163
$\text{FFNet}_{\textit{sr}}$ 91.2 13.7 145.1 133.2 32.69/0.9009 29.01/0.7910 27.84/0.7449 26.93/0.8095 31.73/0.9209
Table 4: Comparisons on representative benchmark datasets for super-resolution networks at $\times$4 scale. Performance metrics are calculated on the Y-channel. All computational statistics are calculated corresponding to an HR image of size 1280$\times$720. Memory denotes maximum GPU memory consumption during the inference phase. Best and second-best results are bolded and underlined, respectively.
We evaluate the effectiveness of our FFNet on two representative benchmark datasets: ADE20K [120] and Cityscapes [17]. For backbone-level comparison, our models are trained for 160k-iterations with a batch size of 16, following standard recipe [58, 59]. For system-level comparisons, we adopt the same training settings in SegNeXt [34] and use ImageNet-1k pretrained models for initialization. We also use MMSegmentation toolbox [15] for all semantic segmentation experiments.
Results. Tab. 3 shows that our methods excel not only as backbones but also as standalone segmentors, highlighting our model’s superior performance-latency trade-off. For instance, FFNet-4 achieves a $0.6$ mIoU improvement over MogaNet-B while being 1.9$\times$/2.5$\times$ faster on the GPU/mobile device, respectively. In addition, $\text{FFNet}_{\textit{seg}}$ consistently outperforms task-specific prior arts.
3D Sem. Seg. ScanNet [18] ScanNet200 [77] S3DIS [1]
Model Val Test Val Test Area5 6-fold
PointNeXt [72] 71.5 71.2 - - 70.5 74.9
ST [46] 74.3 73.7 - - 72.0 -
OctFormer [97] 75.7 76.6 32.6 32.6 - -
Swin3D [108] 76.4 - - - 72.5 76.9
OA-CNNs [70] 76.1 75.6 32.3 33.3 71.1 -
PTv2 [102] 75.4 74.2 30.2 - 71.6 73.5
$\text{FFNet}_{\textit{3D}}$ 76.1 76.5 33.8 34.5 72.6 77.1
Table 5: 3D semantic segmentation.
4.2 Generalizing FFNet beyond Recognition
Efficient Super-Resolution. In this section, we explore the effectiveness of FFNet beyond recognition, extending its evaluation to low-level vision tasks. $\text{FFNet}_{\textit{sr}}$ consists of three parts: shallow feature extractor (HSF), deep feature extractor (HDF), and upscale module (HRec) following residual-in-residual architecture [114]. HSF extracts shallow features from a Low-Resolution (LR) image using a 3$\times$3 convolution. The deep feature extractor HDF, comprising a stack of FFNet blocks, refines shallow features iteratively. We only use 3$\times$3 depthwise convolutions for spatial modeling, with 96 base channels and an FFN expansion ratio of 2. Specifically, $\text{FFNet}_{\textit{sr}}$-light and $\text{FFNet}_{\textit{sr}}$ employ 36 and 48 blocks, respectively. We employ an anchor-based long residual connection [28, 124] where the LR image is repeated RGB-wise and added at the end of the HDF. Finally, HRec reconstructs the High-Resolution (HR) image by up-sampling refined features utilizing pixelshuffle.
$\text{FFNet}_{\textit{sr}}$-light and $\text{FFNet}_{\textit{sr}}$ are trained on the DIV2K [52] and DF2K (DIV2K + Flicker2K [88]) dataset, respectively. Five standard datasets are adopted for testing: Set5 [3], Set14 [111], BSD100 [62], Urban100 [40], and Manga109 [63]. We use PSNR and SSIM to evaluate the SR performance on the Y channel from the YCbCr color space after cropping 4 pixels at boundary [42]. Our models are trained by minimizing L1 loss and frequency reconstruction loss [92, 84, 118] with a weight factor set of 0.05. For more details on the implementation, please refer to Appendix F.2.
Results. In Tab. 4, we compare our models against recent state-of-the-art SR models. Notably, $\text{FFNet}_{\textit{sr}}$ significantly outperforms SR specialists in terms of efficiency, running seamlessly even on resource-constrained devices, demonstrating its feasibility for real-world applications. $\text{FFNet}_{\textit{sr}}$, employing 3$\times$3 depthwise convolutions within the MetaMixer framework, not only captures high-frequency details effectively but also allows for the stacking of more blocks within the same computational budget compared to the baselines [84, 12, 96] that use complex mixers, thereby enhancing its ability to capture long-range contexts. The impressive results suggest that our mixer design is not only effective for high-level tasks such as recognition but also well-suited for solving low-level problems. Visual results of the $\text{FFNet}_{\textit{sr}}$ are provided in Appendix G.1.
4.3 Generalizing FFNet beyond Image
In this section, we introduce models that adapt FFNet for point cloud and time series data by simply substituting its 2D convolutions with 1D/3D convolutions.
@@ 4.3 Generalizing FFNet beyond Image
3D Semantic Segmentation. Following [102], we adopt the U-Net [76] structure, comprising four stages of encoders and decoders with block depths of [2, 2, 6, 2] and [1, 1, 1, 1], respectively. We utilize Grid Pooling [102], setting all grid size multipliers to $\times$2, which represents the expansion ratio over the preceding pooling stage. For token mixing, we employ submanifold convolution [31]. Additionally, to manage memory-constrained scenarios requiring small batch sizes, we use both batch normalization and pre-norm (Layer Norm) scheme. Due to the implementation limitations of submanifold convolution, which does not support depthwise operation, we restrict the kernel size to 3$\times$3$\times$3 to handle overfitting issues. Despite these constraints, the obtained results are promising, demonstrating that even small kernels are highly effective within the MetaMixer framework and U-Net architecture.
We conduct experiments using our $\text{FFNet}_{\textit{3D}}$ on the representative datasets: ScanNetv2 [18], ScanNet200 [77], and S3DIS [1]. We use Pointcep [16] codebase specialized for point cloud perception tasks. Detailed descriptions of architecture and training settings can be found in Appendix F.3.
Results. Our $\text{FFNet}_{\textit{3D}}$ demonstrates superior performance compared to the task-specialized competitors, as shown in Tab. 5. Notably, our architecture design is embarrassingly simple compared to the baselines, reinforcing the idea that the MetaMixer framework is inherently a key performance driver. Moreover, we expect that incorporating adaptive convolution [70] or large kernel convolution [11] into our mixer design could yield further improvements.
@@
Table 6: Multivariate time series long-term forecasting results with prediction lengths $S\in\{96,192,336,720\}$ and fixed lookback length $T=96$. Results are averaged from all prediction lengths. ETT results further averaged by subsets. The baseline numbers are borrowed from iTransformer [57].
[…]
U-DiT-B [87] 22.2G 276.5 3.10 241.56
U-FFNet-B 19.6G 331.0 2.99 245.71
Table 7: ImageNet 256$\times$256 generation performance with classifier-free guidance ($cfg=1.5$). Throughput is measured only for the U-Net, using an NVIDIA RTX A6000.
@@
Time Series Forecasting. Our $\text{FFNet}_{\textit{1D}}$ employs a backbone-head structure with RevIN [43], where the backbone consists of stacked FFNet blocks. In our mixer design, we utilize a 51-sized kernel for 1D depthwise convolution to effectively catch up the global receptive fields of transformer- and MLP-based models. In fact, directly applying the FFNet block (Fig. 3 (c)) to time series data is challenging due to the presence of variable dimension in addition to channel and temporal dimensions in multivariate setting. To effectively capture both temporal and cross-variable dependencies, we modify FFN part using appropriate grouping and feature reshaping. Detailed descriptions of these time series-specific modifications and training settings are available in Appendix F.4.
@@
We conduct long-term forecasting experiments on popular 8 real-world datasets, including ETT (4 subsets), ECL, Exchange, Traffic, Weather. Following [57], we set the input length as 96 for all the results. We compute the Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics for multivariate time series forecasting.
@@
Results. Tab. 7 shows the competitive performance of $\text{FFNet}_{\textit{1D}}$ in multivariate time-series forecasting. Remarkably, $\text{FFNet}_{\textit{1D}}$ achieves overall superior performance compared to the state-of-the-art transformer models, demonstrating the excellent task-generality of our MetaMixer-based convolutional mixer design.
In summary, FFNets not only achieve strong performance but also substantially enhance efficiency across various devices, covering both high-level recognition and low-level vision tasks. Moreover, their seamless extension beyond image data underscores the MetaMixer framework as a crucial architectural foundation.
4.4 Unified Mixer Design with MetaMixer
Our MetaMixer framework, when combined with convolution, provides an efficient baseline while also opening a unified perspective that encompasses attention (see Appendix B and C). While convolutional interaction, such as FFNified attention, is more efficient than attention in high-resolution settings, the advantages of attention remain compelling—it excels in capturing high-level semantic representations and enables flexible operations through inter-token manner. Furthermore, it is noted in [68] that convolution and attention complement each other, as convolution is adept at capturing high-frequency signals while attention specializes in low-frequency signals. This implies that leveraging both worlds within the MetaMixer framework in a balanced manner can optimize both computational efficiency and representational power.
ImageNet Generation. To demonstrate the efficacy of the aforementioned approach, we conduct ImageNet-256 class-conditional generation experiments. Specifically, we choose U-DiT [87] used in Latent Diffusion [75] as our baseline and replace attention with FFNified attention in the encoder/decoder, except for the bottleneck stage. FFNified attention efficiently extracts fine-grained details in high-resolution stages, while attention in the bottleneck stage models semantic information. As shown in Tab. 7, this synergy enables our model to achieve improved FID-Throughput performance. Fig. 14 shows selected samples generated using our ImageNet-256 model.
5 Discussion and Conclusion
Although convolution has been extensively explored before, we are the first to integrate it into query-key-value interactions, introducing a versatile convolutional mixer design that offers general efficiency across various vision tasks and devices. For a more detailed discussion on our relationship with previous CNNs, please refer to Appendix C. Furthermore, by making effective use of large kernels, FFNified attention achieves an effective receptive field comparable to global self-attention (see Appendix D).
@@ 5 Discussion and Conclusion
In this work, we analyzed the FFN from a seldom-explored perspective and integrated self-attention and FFN-based mixers within a query-key-value framework. We further abstracted this framework into a general mixer architecture, termed MetaMixer, where sub-operations are unspecified. Based on the MetaMixer, we presented a convolutional mixer design with FFNified attention. In numerous tasks where transformer-based and convolution-based models are in fierce competition, we merge both approaches in a completely new direction, and achieve a new state of the art with significant efficiency gains. Moreover, the proposed unified mixer design provides deeper insights into the synergy between the two approaches. These achievements underscore that the significance lies not in specific module per se, but in satisfying task-specific functionalities, starting with MetaMixer as the foundation.
[…]
Supplementary Material
[…]
@@ Appendix A FFNs Are Key-Value Memories
We also find that $K_{548}$ and $K_{1461}$ from the smaller variant PoolFormer-S12, as well as $K_{382}$ and $K_{2869}$ from the ViT-B trained on ImageNet-21K, operate similarly to the special keys discussed above. Additionally, Li et al. [51] found that the activations of the FFN in ViT models exhibit sparsity, which aligns with the observations in Fig. 2 (a). The above results suggest that FFNs generally operate as key-value memories across various models.
@@ Appendix B MetaMixer Generality
As shown in Fig. 8, the most commonly used mixers are instantiated by specifying suboperations of the MetaMixer, underscoring that the query-key-value mechanism is a essential requirement for competitive performance. Specifically, the spatial MLP [^1] Both MLP (Multi-Layer Perceptron) and FFN (Feed-Forward Network) generally refer to modules containing two fully-connected layers with an activation function between them. In this paper, we refer to this module as FFN overall, but use MLP specifically in the context of MLP-Mixer [89]. in MLP-Mixer [89] operates via the same query-key-value mechanism as an FFN, but applies it along different axes—token and channel, respectively. The ConvNeXt [59] block can be interpreted as employing a 7$\times$7 depthwise convolution as the query projection layer within an FFN. Additionally, self-attention and FFNified attention utilize inter-token matrix multiplication and depthwise convolution, respectively, to compute query-key-value interactions.
@@ Appendix C Detailed Comparisons with Related Work
MLP-Mixer. The term “FFNification” might initially evoke thoughts of spatial MLPs [89] rather than our proposed FFNified attention due to its direct implications. However, our use of the term encompasses not only the adaptation of mechanisms but also aims to achieve the generality and efficiency characteristic of FFNs. Spatial MLPs, on the other hand, suffer from parameter inefficiency [89, 117] and are impractical for high-resolution inputs, similar to the global self-attention. They are also unable to handle multiple resolutions, limiting their applicability to downstream tasks. Consequently, depthwise convolution emerges as an efficient and effective alternative, addressing these issues. Unlike spatial MLPs, which share the same kernel across all channels, depthwise convolution applies different kernels to each channel. To bridge the receptive field gap between convolution and spatial MLP, we employ large kernels (e.g., 7$\times$7 or 9$\times$9 for image recognition and 51 for time series).
Model Type Top-1 GPU Throughput (imgs/sec) CPU Throughput (imgs/sec) Mobile Macbook Air
(%)     bs 1     bs 64     bs 1     bs 16 Latency (ms) Latency (ms)
MogaNet-T [48] CNN 80.0     463     2169     25     72 2.5 2.0
FFNet-1 CNN 81.3     1430     3965     38     62 1.8 1.3
Swin-T [58] ViT 81.3     -     2324     30     35 24.8 6.9
BiFormer-T [125] ViT 81.4     333     548     4     5 - -
SMT-T [54] ViT 82.2     786     2362     22     25 23.2 15.9
RMT-T [29] ViT 82.4     399     1569     -     - - -
DaViT-T [23] ViT 82.8     312     2120     23     25 12.7 8.3
ConvNeXt-T [59] CNN 82.1     952     2633     41     50 3.7 2.4
FFNet-2 CNN 82.9     776     2163     31     52 3.1 2.1
iFormer-S [81] ViT 83.4     500     1418     23     34 27.3 12.8
CMT-S [33] ViT 83.5     407     1310     13     23 11.3 8.1
Swin-B [58] ViT 83.5     -     841     13     14 33.5 12.1
MaxViT-T [93] ViT 83.6     -     894     17     21 120.1 49.9
GC ViT-T2 [35] ViT 83.7     -     -     18     25 201.4 69.1
SMT-S [54] ViT 83.7     421     1289     12     13 41.5 28.3
BiFormer-S [125] ViT 83.8     162     257     2     3 - -
MogaNet-S [48] CNN 83.4     417     1300     21     47 4.3 3.1
HorNet-S7×7 [74] CNN 83.8     351     954     12     18 5.4 3.7
ConvNeXt-B [59] CNN 83.8     399     1263     16     17 8.1 5.6
FocalNet-B (LRF) [107] CNN 83.9     287     828     12     31 10.0 7.0
FFNet-3 CNN 83.9     490     1431     22     40 4.5 2.9
CMT-Br256 [33] ViT 84.5     294     607     6     10 13.2 9.9
iFormer-Sr384 [81] ViT 84.6     334     403     9     10 80.3 34.8
FFNet-3r384 CNN 84.5     387     595     12     20 9.1 7.1
MaxViT-Tr384 [93] ViT 85.2     149     314     6     6 247.6 91.6
NFNet-F2r384 [4] CNN 85.1     194     347     6     10 27.0 18.0
ConvNeXt-Br384 [59] CNN 85.1     229     268     5     6 21.3 15.4
HorNet-B7×7,r384 [74] CNN 85.3     175     233     5     6 21.3 16.3
FFNet-4r384 CNN 85.3     282     394     10     13 15.2 10.8
@@
Table 8: Comparison of inference speeds across various hardware platforms. “bs” denotes batch size. For CPU benchmarks, models are converted to ONNX runtime format and evaluated on an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz processor using a single thread. For Mobile and M2Air, we export the models using CoreML tools and report the median latency over 1,000 runs.
ConvNeXt. Both the ConvNeXt [59] block and FFNified attention mixer share basic elements such as convolution and GELU activation, yet they differ in the operating axes in the query-key-value mechanism. FFNified attention actively extracts spatial features through two consecutive large-kernel convolutions, calculating query-key and coefficient-value interactions. In contrast, the ConvNeXt block uses depthwise convolution as a means of query projection, insufficiently utilizing the surrounding context during channel interaction. Hence, employing our proposed FFNified attention and ConvNeXt block as token mixer and channel mixer, respectively, provides a compact mixer design space. In FFNified attention, it is natural to not expand channels, leading our FFNet to maintain a larger base channel size compared to similar-sized baseline models and thus reduce the expansion ratio in the channel mixer to 3. This decision aligns with prior research findings [56, 8] that the expanded features of channel mixer contain substantial redundancy. Thanks to our comprehensive design strategy and compact mixer space, our models demonstrate a better performance-speed trade-off than the ConvNeXt models especially on resource-constrained devices. (see Tab. 8).
Vision Transformer (ViT). Self-attention operates at the token level, employing matrix multiplication to calculate inter-token similarity and aggregate values using the resulting coefficients. In contrast, FFNified attention operates at the channel level, detecting local patterns through depthwise convolution and applying a subsequent convolution to the resulting coefficients to aggregate values. To reduce parameter complexity, self-attention shares attention weights across channels, while FFNified attention shares them across positions. Another key difference lies in the key-value generation process: self-attention utilizes instance-specific key-value pairs through linear projections, while FFNified attention initializes key-values as static weights and updates them on the training set.
Efficient ViTs. Vanilla self-attention, being a token-based module, faces limitations in vision tasks with long token lengths (high-resolution tasks). Swin transformer [58] addresses this by performing self-attention within local windows. However, window attention necessitates complex reshape operations, hindering efficient deployment across various inference platforms [56, 110]. Another line of work introduces sparsity into attention [125, 29, 23, 93] or leveraging token merging layer (e.g., strided convolution or pooling) to shorten sequences before applying attention [33, 81]. Yet, these methods demand meticulous architectural tuning and generally lack efficiency on devices other than GPUs. Conversely, FFNified attention, composed solely of hardware-friendly operations like convolution and GELU, exhibits high speed across diverse devices (see Tab. 8).
Figure 9: Visual explanation of re-parameterizing multi-branches for extracting multi-scale features. For a detailed discussion, please refer to Appendix E.
@@
ConvNeXt-B 2.2% 3.7% 7.9% 64.6%
Swin-B 1.0% 1.4% 3.6% 41.4%
FastViT-MA36 1.6% 3.2% 10.0% 97.5%
DeiT-B 3.7% 8.2% 24.5% 98.3%
FFNet-3 15.5% 23.8% 40.9% 98.6%
@@
Figure 10: Left. Effective receptive field (ERF) comparison. FFNet has broader ERFs compared to Swin Transformer [58] and ConvNeXt [59], despite using the same kernel (window) size. Right. Quantitative analysis on the ERF with the high-contribution area ratio $r$. A higher value of $r$ indicates a smoother distribution of high-contribution pixels, which implies a larger ERF.
CNNs. HorNet [74], FocalNet [107], and MogaNet [48] highlight the importance of multi-order interactions and propose high-order convolutional interaction modules. SMT [54] employs multi-head mixed convolution to efficiently extract multi-scale representations. In contrast, by compactly utilizing convolution within the more fundamental query-key-value framework (i.e. MetaMixer), rather than multi-order or multi-scale interactions, our models achieve an optimal performance-speed trade-off (see Tab. 8). On the other hand, large-kernel CNNs [26, 25, 55] depend on specialized libraries like GEMM, which hinders their general deployment across diverse environments such as TensorRT and CoreML.
Appendix D Effective Receptive Field Analysis